OpenStudy (anonymous):

1. (g + 4) – 3g = 1 + g (1 point

OpenStudy (comm.dan):

What do you think you do as your first step?

OpenStudy (anonymous):

1-3

OpenStudy (anonymous):

Simple..you first +3g both sides so it ends up as (g+4)=1+4g and then you subtract 1 from both sides so now it's g+3=4g then subtract a g from both sides and you get 3= 3g. now divide both sides by 3 and you get g=1

OpenStudy (comm.dan):

First, I would distribute then tell me what you get after that Combine the like terms one both sides of the equation.

OpenStudy (anonymous):

OH! Thank you both your were a great help! I got it now :)

OpenStudy (comm.dan):

Any more questions

OpenStudy (comm.dan):

??

OpenStudy (anonymous):

I wont choose the best because both were good. No thanks.

OpenStudy (anonymous):

I was better. lolol jk

OpenStudy (anonymous):

That's not nice :/

OpenStudy (anonymous):

Thank you!

OpenStudy (comm.dan):

No problem